# FastSeg: Efficient Training-Free Open-Vocabulary Segmentation via Hierarchical Attention Refinement Method

### v1

@@ Abstract
Open-vocabulary semantic segmentation (OVSS) aims to segment objects from arbitrary text categories without requiring densely annotated datasets. Although contrastive learning based models enable zero-shot segmentation, they often lose fine spatial precision at pixel level, due to global representation bias. In contrast, diffusion-based models naturally encode fine-grained spatial features via attention mechanisms that capture both global context and local details. However, they often face challenges in balancing the number of iterations with the quality of the segmentation. In this work, we propose FastSeg, a novel and efficient training-free framework with only (1+1)-step of reverse process of a pretrained diffusion model (e.g., Stable Diffusion). Moreover, instead of running multiple times for different classes, FastSeg performs segmentation for all classes at once. To further enhance the segmentation quality, FastSeg introduces three key components: (i) a dual-prompt mechanism for discriminative, class-aware attention extraction, (ii) a Hierarchical Attention Refinement Method (HARD) that enhances fused cross-attention using scale-aligned self-attention maps, and (iii) a Test-Time Flipping (TTF) scheme designed to improve spatial consistency. Extensive experiments show that FastSeg achieves state-of-the-art training-free performance, obtaining 43.8% average mIoU across PASCAL VOC, PASCAL Context, and COCO Object benchmarks while maintaining superior inference efficiency. Our results demonstrate that FastSeg provides a strong foundation for extendability, bridging the gap between segmentation quality and inference efficiency.
@@ 1 Introduction
Figure 1: Open-vocabulary semantic segmentation results using FastSeg. We visualize predictions on challenging scenarios, including contextual images, camouflaged objects, and internet-collected images. The highlighted results demonstrate that, in addition to accurate segmentation of both stuff and things, our method effectively handles camouflaged objects, distinguishes between different characters or dog breeds, and differentiates objects based on color variations.
@@
In this paper, we propose FastSeg, an efficient zero-shot framework for open-vocabulary semantic segmentation (OVSS) that leverages text-to-image diffusion models, as shown in Figure 1. FastSeg balances high segmentation quality with computational efficiency through key innovations. It performs image reconstruction using DDIM Inversion, enabling the simultaneous extraction of both self-attention and cross-attention maps across all candidate classes. A dual-prompt mechanism—a text prompt for reconstruction and a class prompt for generating class-specific cross-attention maps—enables multi-class segmentation without repeated inference. Additionally, FastSeg introduces the Hierarchical Attention Refinement Method (HARD), which aggregates attention maps from multiple scales, combining fine details from higher resolutions and broader context from lower resolutions to improve localization, shape accuracy, and boundary delineation. Finally, the Test-Time Flipping (TTF) strategy enhances segmentation accuracy by averaging predictions from both the original and flipped images, improving spatial consistency and robustness.
@@
    We introduce FastSeg, a novel training-free open-vocabulary segmentation method that leverages diffusion models with fast (1+1)-step DDIM inversion and a unified class prompt, enabling efficient attention map generation for all candidate classes in a single inference run.
@@
    We summarize our key contributions as follows: We conduct comprehensive experiments on three benchmark datasets—PASCAL VOC, PASCAL Context, and COCO Object—demonstrating that FastSeg achieves state-of-the-art performance among training-free methods while maintaining superior computational efficiency.
@@ 2.2 Semantic Segmentation with Diffusion Models
Figure 2: Overview of the proposed FastSeg. he input consists of both the original and horizontally flipped images, which are first reconstructed via DDIM inversion to obtain their latent representations. In parallel, a text prompt $\mathcal{P}$ and a class prompt $\mathcal{P^{\prime}}$ are generated to guide inversion and extract class-specific cross-attention maps corresponding to candidate classes (e.g., bus, motorbike, sheep). Cross-attention and self-attention maps at multiple resolutions are then extracted, transformed, and fused via weighted aggregation, where the self-attention maps are used to refine the class-specific cross-attention maps. Final segmentation masks are derived from the refined per-class score maps. Test-Time Flipping (TTF) is applied during inference to enhance robustness: attention maps from the flipped image are spatially realigned and averaged with those from the original image to improve prediction reliability.
@@ 3 Methods
3.1 FastSeg Overview
Our objective is to perform training-free open-vocabulary semantic segmentation, which assigns each pixel to one of the candidate categories $\mathcal{C}=\{c_{1},c_{2},...,c_{M}\}$ without requiring dense pixel-level annotations. Given a real image $I_{0}$ and a candidate class set $\mathcal{C}$, our proposed FastSeg framework extracts segmentation masks via cross-attention and self-attention analysis from a pre-trained Stable Diffusion (SD) model. The overall framework is depicted in Figure 2. The entire pipeline consists of three stages:
@@ 3.2 Prompt Construction
To enable semantic reconstruction and class-aware attention extraction, FastSeg introduces a dual-prompt mechanism composed of a textual prompt $\mathcal{P}$ and a class prompt $\mathcal{P}^{\prime}$. The textual prompt $\mathcal{P}$, generated by an image captioning model (e.g., BLIP), serves as the text condition for image reconstruction via DDIM inversion, facilitating faithful image generation. To explicitly inform the model about candidate classes, we append the class prompt $\mathcal{P}$ to the textual prompt, following the structure: "image caption; $\mathcal{P}$’". This design enables the diffusion model to focus on class-specific semantics during generation. The class prompt $\mathcal{P}^{\prime}$ is constructed from a candidate label set $\mathcal{C}=\{c_{1},\ldots,c_{M}\}$, containing of $M$ open-vocabulary categories. For example, if $\mathcal{C}$ = { tv monitor, dog, person }, then $\mathcal{P}^{\prime}=``tv{\color[rgb]{1,1,1}s}monitor,$ $dog,$ $person"$. These candidate labels are essential for computing class-specific cross-attention maps that connect textual concepts to spatial regions in the semantic reconstruction of the image. Unlike contrastive models such as CLIP, which inherently suppress irrelevant concepts while enhancing relevant ones, text-to-image models like Stable Diffusion (SD) do not automatically ignore out-of-context objects. Instead, SD can preserve unrelated visual concepts as semantic noise. This behavior is illustrated in Figure 3, where distractor classes (those not present in the image) still produce activated attention maps. This motivates the need for a tailored class candidate set $\mathcal{C}$ for each image.
@@ 3.3.3 Attention Extraction for Segmentation
FastSeg generates semantic segmentation masks by extracting cross-attention and self-attention maps during the Stable Diffusion denoising process (latent $z_{1}$ to $z_{0}$). The cross-attention maps $\mathcal{C}^{l,r}\in[0,1]^{r\times r\times d_{\mathcal{P}}}$ capture spatial correlations between text tokens and image regions. On the other hand, self-attention maps $\mathcal{S}^{l,r}\in[0,1]^{r\times r\times r\times r}$ preserve the spatial structure of the image by modeling pairwise similarities between image positions. These attention maps are computed across multiple resolutions =$\mathcal{R}=\{8,16,32,64\}$, with each resolution having $L$ layers. The maps for each resolution are averaged over layers to obtain a final representation:
@@ 4 Experiments
Table 1: Comparison with existing methods. Models in the first three rows are finetuned on target datasets while the rest approaches do not require mask annotations. Bold fonts refer to the best results and underline fonts refer to the second best.
@@ 4.0.1 Datasets and Metric
We evaluate the proposed FastSeg on standard benchmark datasets for the Open-Vocabulary Semantic Segmentation (OVSS) task: PASCAL VOC, PASCAL Context, and MS COCO Object, comprising 20, 59, and 80 object categories, respectively (excluding the background class). As FastSeg is a training-free method, evaluation is conducted directly on the validation sets of each dataset, which include 1,449 images from PASCAL VOC, 5,004 images from PASCAL Context, and 5,000 images from MS COCO Object. To assess segmentation performance and enable fair comparison with existing approaches, we employ the mean Intersection-over-Union (mIoU) metric, which is a widely used standard for evaluating semantic segmentation models that measure the overlap between predicted and ground-truth regions across classes.
@@ 4.0.2 Implementation Details
We employ 2-Rectified Flow, a distilled variant of Stable Diffusion 1.5 (SD 1.5), as the primary model for attention map extraction. All inferences of SD are performed at an image resolution of 512$\times$512 pixels. In the multi-resolution attention fusion module, we assign distinct weights to cross-attention and self-attention maps. Specifically, for cross-attention maps, we use the weight vector $\{0.15,0.7,0.15,0\}$ corresponding to resolution levels $\{8,16,32,64\}$, while self-attention maps, we assign weights of $\{0.1,0.1,0.5,0.3\}$. These weights are determined through both qualitative inspection of attention heatmaps and quantitative evaluation of segmentation performance at each resolution (see Table 5). Rather than relying on extensive hyperparameter tuning, we prioritize semantic relevance, assigning greater weights to layers that contribute more effectively to accurate segmentation. All experiments are implemented in PyTorch and conducted on a system equipped with an NVIDIA GeForce RTX A5000 GPU (32GB RAM) and an Intel(R) Core(TM) i9-10900X CPU. FastSeg is executed using mixed precision (16-bit floating point), requiring approximately 13.4 GB of GPU memory with a batch size of 2, and achieves an average inference time of 0.36 seconds per image (excluding candidate class extraction).
@@ 4.1.1 Quantitative results
Table 1 summarizes the performance of our proposed FastSeg method across three widely-used open-vocabulary semantic segmentation benchmarks: PASCAL VOC 2012, PASCAL Context, and COCO Object. FastSeg achieves mIoU scores of 64.0% on PASCAL VOC, 31.3% on PASCAL Context, and 36.2% on COCO Object, leading to the highest average mIoU among training-free diffusion-based approaches across all three datasets. While results from contrastive vision-language models are also reported for comprehensive comparison, our primary focus centers on a direct and fair comparison with recent training-free diffusion-based methods, including Diffusion Baseline [39], FreeSeg-Diff [19], DiffCut [38], DiffSeg [36], DiffSegmenter [20], and InvSeg [21]. The performance of Diffusion Baseline is referenced from the InvSeg study [21], while the results of Semantic DiffSeg, an adaptation of DiffSeg [36] for open-vocabulary settings, are taken from its corresponding publication [20]. FastSeg establishes state-of-the-art accuracy on PASCAL VOC 2012 and delivers the highest overall average performance across all datasets. On PASCAL Context and COCO Object, FastSeg ranks second, with results only marginally trailing behind the best-performing methods on each dataset.
@@
3 InvSeg H100 32.4G 7.9s 42.4
DiffSegmentor RTX 4090 16.1G M×$\times$0.44s 41.8
FastSeg (Our) RTX 4090 13.4G 0.36s 43.8
Figure 8: Open vocabulary semantic segmentation results of FastSeg and DiffSegmenter [20], DiffCut [38], ClearCLIP [5], CLIP-DIY [3], OVSegmentor [12]
Table 3: Ablation study on the contributions of each component in FastSeg, including self-attention refinement, cross-attention weighted aggregation, self-attention weighted aggregation, class prompt design, and test-time flipping. Removing any individual component leads to decreased performance.
FastSeg PASCAL VOC PASCAL Context COCO Object $\downarrow\Delta$
@@
Table 2 presents a quantitative evaluation of the operational efficiency among existing state-of-the-art diffusion-based segmentation methods, including DiffSegmentor, InvSeg, and our FastSeg. In terms of memory consumption, FastSeg demonstrates superior efficiency, utilizing only 13.4GB, significantly less than DiffSegmentor’s 16.1GB and InvSeg’s 32.4GB . Notably, regarding inference time, FastSeg achieves superior performance with just 0.36 seconds on an RTX 4090 GPU. This figure represents a substantial improvement over DiffSegmentor, whose inference time depends on the number of candidate classes M, and is particularly impressive when compared to InvSeg, which requires up to 7.9 seconds per image on an NVIDIA H100 GPU due to its complex prompt embedding optimization procedure. These results demonstrate that FastSeg offers high inference efficiency while significantly reducing the necessary computational resources.
In summary, FastSeg not only achieves state-of-the-art segmentation performance among training-free diffusion-based methods but also demonstrates superior inference efficiency and deployment simplicity, making it an efficient and scalable solution for real-world open-vocabulary semantic segmentation.
@@ 4.1.2 Qualitative results
Figure 8 provides a qualitative comparison of FastSeg against several methods, including DiffSegmenter, DiffCut, ClearCLIP, CLIP-DIY, and OVSegmentor, on two benchmark datasets: PASCAL VOC 2012 (top 7 rows) and PASCAL Context (bottom 6 rows). To ensure fairness, the PAMR post-processing step [46] was removed when showing DiffCut results. The results highlight several key observations. DiffSegmenter performs well in capturing object shapes on PASCAL VOC but struggles with contextual classes (e.g., tree, glass) in PASCAL Context. Additionally, its performance is highly sensitive to the quality of the candidate class list. DiffCut produces unstable segmentation masks, attributed to its reliance on an unsupervised grouping strategy and requiring FC-CLIP for semantic label association, making it suboptimal for open-vocabulary segmentation tasks. Among CLIP-based methods, CLIP-DIY often suffers from false positives and false negatives, while ClearCLIP partially mitigates this issue but still shows failure cases—particularly missing detections for classes like person. This indicates limited spatial localization capacity in these models. OVSegmentor exhibits strong geometric segmentation but lacks effective mechanisms to filter out irrelevant or negative classes, leading to noisy outputs and reduced precision.
In contrast, FastSeg consistently delivers high-quality segmentation masks, demonstrating both accurate geometric localization and precise category alignment, even in challenging scenarios. These results substantiate FastSeg’s advantage in open-vocabulary settings, offering a strong balance between semantic accuracy and spatial precision.

### v2

# SwiftSeg: Efficient Training-Free Open-Vocabulary Segmentation via Hierarchical Attention Refinement Method

@@ Abstract
Open-vocabulary semantic segmentation (OVSS) aims to segment objects from arbitrary text categories without requiring densely annotated datasets. Although contrastive learning based models enable zero-shot segmentation, they often lose fine spatial precision at pixel level, due to global representation bias. In contrast, diffusion-based models naturally encode fine-grained spatial features via attention mechanisms that capture both global context and local details. However, they often face challenges in balancing the number of iterations with the quality of the segmentation. In this work, we propose SwiftSeg, a novel and efficient training-free framework with only (1+1)-step of reverse process of a pretrained diffusion model. Moreover, instead of running multiple times for different classes, SwiftSeg performs segmentation for all classes at once. To further enhance the segmentation quality, SwiftSeg introduces three key components: (i) a dual-prompt mechanism for discriminative, class-aware attention extraction, (ii) a Hierarchical Attention Refinement Method (HARD) that enhances fused cross-attention using scale-aligned self-attention maps, and (iii) a Test-Time Flipping (TTF) scheme designed to improve spatial consistency. Extensive experiments show that SwiftSeg achieves state-of-the-art training-free performance, obtaining 43.8% average mIoU across PASCAL VOC, PASCAL Context, and COCO Object benchmarks while maintaining superior inference efficiency. Our results demonstrate that SwiftSeg provides a strong foundation for extendability, bridging the gap between segmentation quality and inference efficiency.
@@ 1 Introduction
Figure 1: Open-vocabulary semantic segmentation results using SwiftSeg. We visualize predictions on challenging scenarios, including contextual images, camouflaged objects, and internet-collected images. The highlighted results demonstrate that, in addition to accurate segmentation of both stuff and things, our method effectively handles camouflaged objects, distinguishes between different characters or dog breeds, and differentiates objects based on color variations.
@@
In this paper, we propose SwiftSeg, an efficient zero-shot framework for open-vocabulary semantic segmentation (OVSS) that leverages text-to-image diffusion models, as shown in Figure 1. SwiftSeg balances high segmentation quality with computational efficiency through key innovations. It performs image reconstruction using DDIM Inversion, enabling the simultaneous extraction of both self-attention and cross-attention maps across all candidate classes. A dual-prompt mechanism—a text prompt for reconstruction and a class prompt for generating class-specific cross-attention maps—enables multi-class segmentation without repeated inference. Additionally, SwiftSeg introduces the Hierarchical Attention Refinement Method (HARD), which aggregates attention maps from multiple scales, combining fine details from higher resolutions and broader context from lower resolutions to improve localization, shape accuracy, and boundary delineation. Finally, the Test-Time Flipping (TTF) strategy enhances segmentation accuracy by averaging predictions from both the original and flipped images, improving spatial consistency and robustness.
@@
    We introduce SwiftSeg, a novel training-free open-vocabulary segmentation method that leverages diffusion models with fast (1+1)-step DDIM inversion and a unified class prompt, enabling efficient attention map generation for all candidate classes in a single inference run.
@@
    We summarize our key contributions as follows: We conduct comprehensive experiments on three benchmark datasets—PASCAL VOC, PASCAL Context, and COCO Object—demonstrating that SwiftSeg achieves state-of-the-art performance among training-free methods while maintaining superior computational efficiency.
@@ 2.2 Semantic Segmentation with Diffusion Models
Figure 2: Overview of the proposed SwiftSeg. he input consists of both the original and horizontally flipped images, which are first reconstructed via DDIM inversion to obtain their latent representations. In parallel, a text prompt $\mathcal{P}$ and a class prompt $\mathcal{P^{\prime}}$ are generated to guide inversion and extract class-specific cross-attention maps corresponding to candidate classes (e.g., bus, motorbike, sheep). Cross-attention and self-attention maps at multiple resolutions are then extracted, transformed, and fused via weighted aggregation, where the self-attention maps are used to refine the class-specific cross-attention maps. Final segmentation masks are derived from the refined per-class score maps. Test-Time Flipping (TTF) is applied during inference to enhance robustness: attention maps from the flipped image are spatially realigned and averaged with those from the original image to improve prediction reliability.
@@ 3 Methods
3.1 SwiftSeg Overview
Our objective is to perform training-free open-vocabulary semantic segmentation, which assigns each pixel to one of the candidate categories $\mathcal{C}=\{c_{1},c_{2},...,c_{M}\}$ without requiring dense pixel-level annotations. Given a real image $I_{0}$ and a candidate class set $\mathcal{C}$, our proposed SwiftSeg framework extracts segmentation masks via cross-attention and self-attention analysis from a pre-trained Stable Diffusion (SD) model. The overall framework is depicted in Figure 2. The entire pipeline consists of three stages:
@@ 3.2 Prompt Construction
To enable semantic reconstruction and class-aware attention extraction, SwiftSeg introduces a dual-prompt mechanism composed of a textual prompt $\mathcal{P}$ and a class prompt $\mathcal{P}^{\prime}$. The textual prompt $\mathcal{P}$, generated by an image captioning model (e.g., BLIP), serves as the text condition for image reconstruction via DDIM inversion, facilitating faithful image generation. To explicitly inform the model about candidate classes, we append the class prompt $\mathcal{P}$ to the textual prompt, following the structure: "image caption; $\mathcal{P}$’". This design enables the diffusion model to focus on class-specific semantics during generation. The class prompt $\mathcal{P}^{\prime}$ is constructed from a candidate label set $\mathcal{C}=\{c_{1},\ldots,c_{M}\}$, containing of $M$ open-vocabulary categories. For example, if $\mathcal{C}$ = { tv monitor, dog, person }, then $\mathcal{P}^{\prime}=``tv{\color[rgb]{1,1,1}s}monitor,$ $dog,$ $person"$. These candidate labels are essential for computing class-specific cross-attention maps that connect textual concepts to spatial regions in the semantic reconstruction of the image. Unlike contrastive models such as CLIP, which inherently suppress irrelevant concepts while enhancing relevant ones, text-to-image models like Stable Diffusion (SD) do not automatically ignore out-of-context objects. Instead, SD can preserve unrelated visual concepts as semantic noise. This behavior is illustrated in Figure 3, where distractor classes (those not present in the image) still produce activated attention maps. This motivates the need for a tailored class candidate set $\mathcal{C}$ for each image.
@@ 3.3.3 Attention Extraction for Segmentation
SwiftSeg generates semantic segmentation masks by extracting cross-attention and self-attention maps during the Stable Diffusion denoising process (latent $z_{1}$ to $z_{0}$). The cross-attention maps $\mathcal{C}^{l,r}\in[0,1]^{r\times r\times d_{\mathcal{P}}}$ capture spatial correlations between text tokens and image regions. On the other hand, self-attention maps $\mathcal{S}^{l,r}\in[0,1]^{r\times r\times r\times r}$ preserve the spatial structure of the image by modeling pairwise similarities between image positions. These attention maps are computed across multiple resolutions =$\mathcal{R}=\{8,16,32,64\}$, with each resolution having $L$ layers. The maps for each resolution are averaged over layers to obtain a final representation:
@@ 4 Experiments
Table 1: Comparison with existing methods. Models in the first three rows are finetuned on target datasets while the rest approaches do not require mask annotations. Bold fonts refer to the best results and underline fonts refer to the second best. Note that “✘” refers to the training-free method, and “–” represents unprovided data.
@@ 4.0.1 Datasets and Metric
We evaluate the proposed SwiftSeg on standard benchmark datasets for the Open-Vocabulary Semantic Segmentation (OVSS) task: PASCAL VOC, PASCAL Context, and MS COCO Object, comprising 20, 59, and 80 object categories, respectively (excluding the background class). As SwiftSeg is a training-free method, evaluation is conducted directly on the validation sets of each dataset, which include 1,449 images from PASCAL VOC, 5,004 images from PASCAL Context, and 5,000 images from MS COCO Object. To assess segmentation performance and enable fair comparison with existing approaches, we employ the mean Intersection-over-Union (mIoU) metric, which is a widely used standard for evaluating semantic segmentation models that measure the overlap between predicted and ground-truth regions across classes.
@@ 4.0.2 Implementation Details
We employ 2-Rectified Flow, a distilled variant of Stable Diffusion 1.5 (SD 1.5), as the primary model for attention map extraction. All inferences of SD are performed at an image resolution of 512$\times$512 pixels. In the multi-resolution attention fusion module, we assign distinct weights to cross-attention and self-attention maps. Specifically, for cross-attention maps, we use the weight vector $\{0.15,0.7,0.15,0\}$ corresponding to resolution levels $\{8,16,32,64\}$, while self-attention maps, we assign weights of $\{0.1,0.1,0.5,0.3\}$. These weights are determined through both qualitative inspection of attention heatmaps and quantitative evaluation of segmentation performance at each resolution (see Table 5). Rather than relying on extensive hyperparameter tuning, we prioritize semantic relevance, assigning greater weights to layers that contribute more effectively to accurate segmentation. All experiments are implemented in PyTorch and conducted on a system equipped with an NVIDIA GeForce RTX A5000 GPU (32GB RAM) and an Intel(R) Core(TM) i9-10900X CPU. SwiftSeg is executed using mixed precision (16-bit floating point), requiring approximately 13.4 GB of GPU memory with a batch size of 2, and achieves an average inference time of 0.36 seconds per image (excluding candidate class extraction).
@@ 4.1.1 Quantitative results
Table 1 summarizes the performance of our proposed SwiftSeg method across three widely-used open-vocabulary semantic segmentation benchmarks: PASCAL VOC 2012, PASCAL Context, and COCO Object. SwiftSeg achieves mIoU scores of 64.0% on PASCAL VOC, 31.3% on PASCAL Context, and 36.2% on COCO Object, leading to the highest average mIoU among training-free diffusion-based approaches across all three datasets. While results from contrastive vision-language models are also reported for comprehensive comparison, our primary focus centers on a direct and fair comparison with recent training-free diffusion-based methods, including Diffusion Baseline [39], FreeSeg-Diff [19], DiffCut [38], DiffSeg [36], DiffSegmenter [20], and InvSeg [21]. The performance of Diffusion Baseline is referenced from the InvSeg study [21], while the results of Semantic DiffSeg, an adaptation of DiffSeg [36] for open-vocabulary settings, are taken from its corresponding publication [20]. SwiftSeg establishes state-of-the-art accuracy on PASCAL VOC 2012 and delivers the highest overall average performance across all datasets. On PASCAL Context and COCO Object, SwiftSeg ranks second, with results only marginally trailing behind the best-performing methods on each dataset.
@@
InvSeg [21] H100 32.4G 7.9s 42.4
DiffSegmentor [20] RTX 4090 16.1G M×$\times$0.44s 41.8
SwiftSeg (Our) RTX 4090 13.4G 0.36s 43.8
Figure 8: Open vocabulary semantic segmentation results of SwiftSeg and DiffSegmenter [20], DiffCut [38], ClearCLIP [5], CLIP-DIY [3], OVSegmentor [12]
Table 3: Ablation study on the contributions of each component in SwiftSeg, including self-attention refinement, cross-attention weighted aggregation, self-attention weighted aggregation, class prompt design, and test-time flipping. Removing any individual component leads to decreased performance.
SwiftSeg PASCAL VOC PASCAL Context COCO Object $\downarrow\Delta$
@@
Table 2 presents a quantitative evaluation of the operational efficiency among existing state-of-the-art diffusion-based segmentation methods, including DiffSegmentor, InvSeg, and our SwiftSeg. In terms of memory consumption, SwiftSeg demonstrates superior efficiency, utilizing only 13.4GB, significantly less than DiffSegmentor’s 16.1GB and InvSeg’s 32.4GB . Notably, regarding inference time, SwiftSeg achieves superior performance with just 0.36 seconds on an RTX 4090 GPU. This figure represents a substantial improvement over DiffSegmentor, whose inference time depends on the number of candidate classes M, and is particularly impressive when compared to InvSeg, which requires up to 7.9 seconds per image on an NVIDIA H100 GPU due to its complex prompt embedding optimization procedure. These results demonstrate that SwiftSeg offers high inference efficiency while significantly reducing the necessary computational resources.
In summary, SwiftSeg not only achieves state-of-the-art segmentation performance among training-free diffusion-based methods but also demonstrates superior inference efficiency and deployment simplicity, making it an efficient and scalable solution for real-world open-vocabulary semantic segmentation.
@@ 4.1.2 Qualitative results
Figure 8 provides a qualitative comparison of SwiftSeg against several methods, including DiffSegmenter, DiffCut, ClearCLIP, CLIP-DIY, and OVSegmentor, on two benchmark datasets: PASCAL VOC 2012 (top 7 rows) and PASCAL Context (bottom 6 rows). To ensure fairness, the PAMR post-processing step [46] was removed when showing DiffCut results. The results highlight several key observations. DiffSegmenter performs well in capturing object shapes on PASCAL VOC but struggles with contextual classes (e.g., tree, glass) in PASCAL Context. Additionally, its performance is highly sensitive to the quality of the candidate class list. DiffCut produces unstable segmentation masks, attributed to its reliance on an unsupervised grouping strategy and requiring FC-CLIP for semantic label association, making it suboptimal for open-vocabulary segmentation tasks. Among CLIP-based methods, CLIP-DIY often suffers from false positives and false negatives, while ClearCLIP partially mitigates this issue but still shows failure cases—particularly missing detections for classes like person. This indicates limited spatial localization capacity in these models. OVSegmentor exhibits strong geometric segmentation but lacks effective mechanisms to filter out irrelevant or negative classes, leading to noisy outputs and reduced precision.
In contrast, SwiftSeg consistently delivers high-quality segmentation masks, demonstrating both accurate geometric localization and precise category alignment, even in challenging scenarios. These results substantiate SwiftSeg’s advantage in open-vocabulary settings, offering a strong balance between semantic accuracy and spatial precision.